OpenStudy (hba):

what is the distance in space between the points with coordinates (-3,6,7) and (2,-1,4)

OpenStudy (hba):

dont tell me the answer

OpenStudy (amistre64):

d^2 = x'^2 + y'^2 + z'^2

OpenStudy (amistre64):

where the 's are the difference between component parts

OpenStudy (hba):

25+49+9=83

OpenStudy (hba):

9.11 right

OpenStudy (hba):

ive used distance formula

OpenStudy (amistre64):

think of the distance formula as the sqrt of the difference of components: the distance from 5 to 16 is then:\(\sqrt{(5-16)^2}\) the distance from 5,3 to 16,4 is then:\(\sqrt{(5-16)^2+(3-4)^2}\) the distance from 5,3,z1 to 16,4,z2 is then:\(\sqrt{(5-16)^2+(3-4)^2+(z_1-z_2)^2}\) etc...